In three years, Herman's grandfather will be six times as old as Herman was last year. When Herman's current age is added to his grandfathers present age, the total is 68. How old are Herman and his grandfather now.
Answer by ikleyn(52802)

.
G + 3 = 6*(H-1)    (1)   ("In three years, Herman's grandfather will be six times as old as Herman was last year. ")

H + G = 68         (2)   ("When Herman's current age is added to his grandfathers present age, the total is 68.")


From (2), express G = 68 - H and substitute it into (1). You will get

(68-H) + 3 = 6*(H-1),

71 - H = 6H - 6,

71 + 6 = 6H + H,

77 = 7H  ====>  H = 11.


Answer.  Herman is 11 years old.  His grandfather is 68-H = 68 - 11 = 57 years old.


-------------------
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.